OpenStudy (anonymous):

Question 6 Assume that in the circuit above, V0=10V, T=1ms, Rw=2Ω, Lw=1mH and RL=3Ω. What is the time constant (in milliseconds) associated with the rise of voltage vR? What is the time constant (in milliseconds) associated with the fall of voltage vR? What is the value of Vmax in Volts? What is the value of Vmin in Volts?

OpenStudy (nurali):

@AmnaAftab Question 6 1. 0.2 2. 0.2 3. 6 4. 0.042

OpenStudy (nurali):

@AmnaAftab If correct please click best response.

OpenStudy (anonymous):

nurali r u sure is this correct answers

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hi guys plz what about those values Assume that in the circuit above, V0=10V, T=0.5ms, Rw=1Ω, Lw=0.5mH and RL=4Ω. plz and thanks

OpenStudy (nurali):

1. 0.1 2. 0.1 3. 7.946 4. 0.0535

OpenStudy (anonymous):

sorry i meand Q4 is those expression correct 1/ Vc(t)=V*(1-exp(-t/R1*c)+exp(-t/R2c)) 2/ Vc(t)=Vc*cos((t-T)/sqrt(L*C)) 3/ iL=-(V/sqrt(L*C))*sin(t-T/sqrt(L*c))

OpenStudy (anonymous):

and thanks in advance nurali

OpenStudy (nurali):

1. V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) 2. (V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) 3. C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C)

OpenStudy (anonymous):

thanks lot nurali :)

OpenStudy (anonymous):

Hi! Nurali need help here.. in q6 My val v=5, rw=3, rl=2, t=0.5m, l=1mh

OpenStudy (anonymous):

Assume that in the circuit above, V0=5V, T=0.5ms, Rw=3Ω, Lw=1mH and RL=2Ω. plz and thanks

OpenStudy (anonymous):

Hi! mjw2 how many tries are left with u??

OpenStudy (anonymous):

I failed two times, What if V0=5V T=0.ms RW=1Ω LW=1mH y RL=1Ω, This is my last question and I need it to get the 60%

OpenStudy (anonymous):

T=0.5 ms

OpenStudy (nurali):

@HEMT 0.2 0.2 5.957 0.042

OpenStudy (anonymous):

Snnnf!!!!!! I got this second x. May data is

OpenStudy (nurali):

@peruana 0.5 0.5 1.811 0.697

OpenStudy (anonymous):

Thanks a lot Naurali, you make my day and save my live.

OpenStudy (anonymous):

I got my 60%, now I'm awaiting my certificate.

OpenStudy (anonymous):

Merry Christmas to all

OpenStudy (nurali):

Same to u.